OpenStudy (anonymous):

Recall that a(b+c) = ab+ac and a(b-c) = ab-ac Simplify if possible: 8 - 6x - 2x

OpenStudy (anonymous):

8-8x alternate form would be -8(x-1) simply x=1

jhonyy9 (jhonyy9):

8-6x-2x = 8 - 8x = 8(1-x)